Calculate the moles of acetic acid, molarity of th
Answer #1

Calculate the moles of NaOH:

Molarity =moles/volume of NaOH

0.096*0.0349 L= Moles of NaOH

Moles of NaOH = 3.3504*10-3 moles

The acetic acid present in the vinegar reacts with NaOH with a stiochiometric ratio of 1.

so moles of NaOH = moles of acetic acid

Therefore, the moles of acetic acid in the vinegar solution are 3.3504*10-3 .

Grams of acetic acid = 60 g/mol * 3.3504*10-3 =0.2010 g

b)

grams of vinegar solution= density of vinegar solution x volume of vinegar solution

=1.01 g/mL *500 mL =505 g

  % of acetic acid = grams of acetic acid/grams of vinegar solution*100 =(0.2010g/505 g)*100=0.0398 %